ganeshie8 (ganeshie8):

Easy yet interesting problem ! Show that \(n^{15}+1\) is composite for all integers \(n\ge 2\)

ganeshie8 (ganeshie8):

found this while attempting this problem https://projecteuler.net/problem=421

OpenStudy (anonymous):

can be proved easy using cauchy inequality and fundamental theorem of algebra

ganeshie8 (ganeshie8):

interesting xD i have a different approach, please show the proof :)

OpenStudy (mathmath333):

let \(n^5=x\) \(\large\tt \begin{align} \color{black}{ n^{15}+1\hspace{.33em}\\~\\ =(n^{5})^3+1\hspace{.33em}\\~\\ =(x)^3+1\hspace{.33em}\\~\\ =(x+1)(x^2-x+1)\hspace{.33em}\\~\\ =(n^5+1)(n^{10}-n^5+1)\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

I hope that can be useful

OpenStudy (thomas5267):

How do you tackle that question though? I don't even know how to begin.

OpenStudy (anonymous):

so does that mean n^3+1 is composite ?

OpenStudy (turingtest):

*

OpenStudy (rational):

we can prove a much stronger statement : \(n^{2k+1}+1\) is composite for all integers \(n\ge 2\) and \(k\ge 1\)

OpenStudy (rational):

\[n^{2k+1}+1 = n^{2k+1} - (-1)^{2k+1} = (n+1)(n^{2k} - n^{2k-1}+\cdots+1)\] for \(n\ge 2\), both the factors above are \(\gt 1\) ending the proof.

OpenStudy (rational):

used this identity for factoring http://gyazo.com/9a998e02beaead2b99413acd559c3267

OpenStudy (anonymous):

this derived from binomial theorem

OpenStudy (anonymous):

so this gives sum of squares composite ,sum of cubes composite ,....ect

OpenStudy (anonymous):

i still think ,then why we can write primes ad difference of squares :P

ganeshie8 (ganeshie8):

sum of squares need not be composite

ganeshie8 (ganeshie8):

4^2 + 1^2 is not composite

OpenStudy (anonymous):

i see but it is when the second term is not 1

ganeshie8 (ganeshie8):

how do you prove that ?

OpenStudy (anonymous):

eh , ok i guess i was wrong

ganeshie8 (ganeshie8):

\(n = a^2 + b^2\) is a prime only when \( n\equiv 1 \pmod 4 \)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

so this proof :- n^2=0,1 mod 4 a^2+b^2=0+0=0 mod 4(invalid since it gives a^2+b^2 even) =0+1=1 mod 4 =1+1=2 mod 4 (invalid since it gives a^2+b^2 even)